Section page offset
s p d
2 21 9

How many page tables are there?

What's the size of each of the page tables? (Assume 4 bytes per entry in the page table)
Answer #1

"Section" part of the information represents total number of tables in the memory and "page" part represents the number of entries in the page table.

Number of page tables ?

A) Section part contains 2 bits, with these 2 bits we can form 22 = 4 tables. Therefore we have 4 page tables in the system.

Size of the each page table ?

A) Page part represents the number of entries in each page table and it contains 21 bits, with these 21 bits we can have 221 entries and it is given that each entry is 4 bytes.

Therefore total page table size = 221 * 4 bytes = 2 * 4 Mbytes = 8 Mbytes.      (220 = 1M)